OpenStudy (ammarah):

Find all zeros of the function given one of the zeros f(x)=15x^3-119x^2-10x+16;8

OpenStudy (ammarah):

Sorry its find all real zeros of the function

OpenStudy (tkhunny):

Have you considered Synthetic Division?

OpenStudy (ammarah):

yes i did it

OpenStudy (ammarah):

I got 15x^3-7x^2-66x+518/x+8

OpenStudy (ammarah):

how do i find all the zeros?

OpenStudy (ammarah):

@Callisto @ganeshie8 @Gravityy.boyy @hartnn

OpenStudy (phi):

if you have a zero, it will divide evenly into the polynomial (no remainder) in other words, for a cubic P(x), you would expect P(x)= (x-a)(x-b)(x-c) (in factored form) if you know c is a root, and divide by (x-c) you get (x-a)(x-b) (but in expanded form)

OpenStudy (ammarah):

i got 8,-6,5 for the zeros is that correct?

OpenStudy (phi):

what did you get after dividing by (x-8) ?

OpenStudy (ammarah):

15x^2+x-2

OpenStudy (ammarah):

and then that factored out to( x+6)(x-5)

OpenStudy (phi):

15x^2+x-2 looks good but (x+6)(x-5) = x^2 + x -30

OpenStudy (ammarah):

how would u factor 15x^2+x-2? I times 15 times -2 and then did the factoring...

OpenStudy (phi):

yes, the "a" coefficient not being 1 is a pain. I do the same thing as you, -2*15= -30 list the pairs of numbers that multiply to give 30 1,30 2,15 3,10 5,6 notice the sign of -30 means "opposite signs" notice the + sign on +x means the bigger number is + test which pair adds to +1: -5+6 = 1 so 5,6 is the pair. Now, because "a" is not 1, I look for the number in the pair that divides into a=15 5 does, and we get 3 that tells me (3x ) is one of the factors. also, divide 3 into 15 to get 4 5 for the other factor (3x ) (5x ) next, we need to get a -5 and a +6, so fill in the blanks. (3x -1) (5x ) (the -1 * 5x gives -5x) (3x -1)(5x +2) (gives +6x)

OpenStudy (phi):

**also, divide 3 into 15 to get 5x for the other factor

OpenStudy (phi):

now solve (3x -1)(5x +2)=0

OpenStudy (ammarah):

ok i get it thanks... i need help on anothquestion...

OpenStudy (ammarah):

factor the polynomial: 2x^3+16y^3

OpenStudy (ammarah):

i know im supposed to use (a+b)(a^2-ab+b^2)

OpenStudy (phi):

the first thought is factor out a 2 \[ 2(x^3 + 8y^3)\] the 2nd thought is 8 = 2^3, so this is the same as \[2( x^3 +(2y)^3)) \]

OpenStudy (phi):

now you can use your expression, with a= x and b = 2y

OpenStudy (ammarah):

so its (x+2y)(x^2-2yx+2y^2)

OpenStudy (phi):

(it's not its) but with a leading 2 ... from 2(x^3 + (2y)^3)

OpenStudy (phi):

and the last term in parens is (2y)^2 = 4y^2 (not 2y^2)

OpenStudy (ammarah):

wait so i multiply the four by each one?

OpenStudy (phi):

(a^3+b^3) factors into (a+b)(a^2-ab+b^2) and 2(a^3+b^3) = 2(a+b)(a^2-ab+b^2) now replace a with x and b with 2y 2(x^3 + 8y^3) =2(x + 2y) ( x^2 -x*2y + (2y)^2 ) or 2x^3 +16y^3= 2(x+2y)(x^2 -2xy +4y^2)

OpenStudy (ammarah):

ohhhi get it could u please help me with polynomial long division?

OpenStudy (phi):

please make it a new post.